the supplement of an angle measures 25 degrees more than twice its complement. Find the measure of the angle.
My work: 180-x= 2(90-x)
180-x= 180-2x
180+x=180
x=1 BUT, x is suppposed to be equal to 25 inthe answer key.
Thank you.

Answer by rapaljer(4671)

I would let x = the complement, since that is at the end of the first sentence.
2x + 25 = the supplement

The equation is based upon the fact that if you subtract the supplement minus the complement, you get 90 degrees.
2x + 25 - x = 90
x + 25 = 90
x= 65 complement
90-x = 25 = the angle

There may be easier ways to do this if we think about it, but at least this works to solve the problem.

OH! I see it!! THE WAY YOU DID IT IS CORRECT!! YOU JUST FORGOT THE 25.
Your equation: 180-x= 2(90-x)

Should have been: 180-x= 2(90-x)+ 25
This works!! x= 25.

R^2 at SCC